QUESTION 25

  1. Let the U.S. dollar-yen spot rate be ¥120/$. Also, let the 180-day forward exchange rate be ¥124.8/$. Then the yen is selling at a per annum _________ of ___________.

a.

premium; 8.00%

b.

premium; 6.30%

c.

discount; 8.00%

d.

discount; 1.57%

QUESTION 26

  1. Assume that a green card holder flies with American Airlines from NYC to Paris. The related transactions will be reported on the Balance of Payments.

True

False
Answer #1

Q 25) Disount; 8%

Reason:- The forward rate for yen in regards to dollar is less than the spot rate, which shows that yen is being sold at discount .

Calculation

Forward discount = [(forward exchange rate - spot exchange rate)/ spot exchange rate]* 360/180

Forward exchange for yen in terms of $/yen = (1/124.8)

= 0.00801282051

Spot exchange rate for yen in terms of $/yen = (1/120)

=0.00833333

Putting the value in above formula

= ( 0.00801282051- 0.00833333333/ 0.00833333333) *360/180

= ( - 0.00032051282/ 0.00833333333) * 2

= ( - 0.0384615384) * 2

= - 0.0769230768. ( Rounding off to 2 decimal)

= - 0.08 or - 8%

Q26 ) True

Reason :- it a payment for service being provided and it will be recorded in current account of balance of payment .